OpenStudy (anonymous):

How do you simplify the square root of 180 and how do you know that it is in simplest form?

OpenStudy (mokeira):

\[\sqrt{180}\] = \[\sqrt{5\times36}\] \[\sqrt{5}\times \sqrt{36}\] \[\sqrt{5}=\sqrt{5}\] \[\sqrt{36} = 6\] so \[6\sqrt{5}\]

OpenStudy (anonymous):

Refer to the attachment from Mathematica v9